The mean of 8 numbers is 11. If an 9th number is now included in the results, the mean becomes 10. What is the value of the 9th number? Produce an algebraic solution.
Answer by edjones(7569)

8*11=88
(88+x)/9=10
88+x=90
x=2
.
Ed